OpenStudy (anonymous):

Find the slope of the line whose equation is 0.05x - 0.03y = 9

OpenStudy (timaashorty):

Slope is 1.6667

OpenStudy (raden):

the slope of line ax + by = c is m = -a/b

OpenStudy (raden):

0.05x - 0.03y = 9 in this case, a = 0.05 and b = -0.03 so, m = -a/b = -0.05/-0.03 = ... ?

OpenStudy (anonymous):

My options are 5/3 3/5 -5/3

OpenStudy (raden):

-0.05/-0.03 = ... ? (it must be in the options)

OpenStudy (anonymous):

5/3?

OpenStudy (raden):

yes, correct

OpenStudy (anonymous):

or -5/3?

OpenStudy (raden):

-0.05/-0.03 = 5/3 -/-, always be +tive